At s.t.p,a certain Mass of gas occupies a volume of 800cm3. Find the temperature at which the gas occupies 1000cm3 and has a pressure of 720mmhg
1
Expert's answer
2021-08-23T02:56:23-0400

P1V1/T1 =P2v2/T2

1 × 800/273= 0.9504 × 1000/T2

2.9304 = 950.4/T2

2.9304 T2 = 950.4

T2 = 324.324K

T2 = 51.32°C